# Models Matter: Setting Accurate Privacy Expectations for Local and Central Differential Privacy

### v1

@@ 4.1.3. Data-Sharing Decision
5. Limitations
Our designs are limited in their focus on a single scenario. Although medical applications are often cited as motivation for studies of DP (Jordon et al., 2018; Bhaskar et al., 2010), DP has not been widely deployed in medical contexts (Dankar and El Emam, 2013). Nevertheless, our privacy labels are transportable to other domains. Future work could transfer our designs to other scenarios and test whether our findings still hold. A limitation of our evaluation is that encountering explanations of DP in practice differs significantly from encountering explanations in an online survey. Future work could investigate comprehension when these explanations are encountered in more natural settings. A third limitation is our focus on a U.S. audience. Our privacy labels may be received differently in a different cultural context. Finally, we present the nature of DP’s protection as binary, when in fact the level of protection depends on the choice of privacy budget. This simplification may be appropriate for small privacy budgets, but the question of determining an acceptable range for the privacy budget is itself a nontrivial problem.
@@ 5. Limitations
6. Discussion
Our results highlight the value of combining disparate best practices from prior work on explaining other security and privacy (S&P) concepts to explain complex PETs such as DP (Kelley et al., 2009; Distler et al., 2020). We find that consequences-focused explanations (i.e., privacy label explanations that highlight information flows) to be a promising approach for promoting accurate understandings of potential data leaks in DP systems. However, to ensure that such explanations are trusted we find that it is necessary to pair such consequences-focused information with a limited amount of high-level information about mechanisms: how DP works to offer particular consequences and protections. Below we discuss potential pitfalls of privacy labels for DP as well as ways to extend our designs to explain other PETs individually or in combination.
@@ 6. Discussion
Although the nutrition label approach shows promise for setting appropriate privacy expectations, it is important to avoid pitfalls from prior deployments of nutrition labels for privacy (Cranor, 2022). For example, iOS privacy labels can be misleading and inaccurate (Kollnig et al., 2022), in part because developers struggle to create accurate labels (Li et al., 2022). Similarly, our labels for DP could be misleading if an organization has implemented DP incorrectly (Jin et al., 2022; Bichsel et al., 2021; Casacuberta et al., 2022; Mironov, 2012; Lyu et al., 2017) or has chosen an inappropriately large privacy budget (Dwork et al., 2019). Specialized programming platforms, audits, and formal verification approaches are therefore an important complement to our work (McSherry, 2009; Reed and Pierce, 2010; Zhang and Kifer, 2017; Tramer et al., 2022; Ding et al., 2018; Kifer et al., 2020), namely by ensuring that the communicated privacy guarantees match the implementation.
Furthermore, while privacy labels can empower individuals to make decisions that better align with their goals and values, it is also important not to overburden individuals in the same way that traditional privacy policies do (McDonald and Cranor, 2008). As some of the participants we interviewed highlighted, it can be difficult to strike the right balance between simplicity and comprehensiveness. Such a balance is important not only for data subjects, but also for other audiences who may encounter DP. For instance, privacy labels for DP could be used to educate policymakers, advocacy organizations, or software developers to support them in various decision-making processes. For example, Mozilla’s “privacy not included” guide offers expert reviews to help buyers choose products that provide strong privacy and security, since it can be difficult for individual buyers to evaluate various data protection policies themselves. One could imagine a similar project to provide reviews for different data collection initiatives. An advocacy organization might use privacy labels for PETs like DP to identify and recommend certain initiatives that provide good S&P guarantees.
Finally, it is crucial that privacy labels for DP be contextual. While the information disclosures our explanations highlight are ones that people care about (Cummings et al., 2021), they represent a starting point which should be used to further adapt explanations for specific contexts. The information disclosures we highlight may not be comprehensive of all specific disclosures people are concerned about across contexts. For example, privacy concerns in a particular educational setting may differ from a medical setting. Future work should also study ways to supplement privacy labels for DP with contextually-appropriate communication about the choice of privacy budget (Nanayakkara et al., 2023; Benthall and Cummings, 2022).
@@
Our approach to designing privacy labels for DP could be adapted to other PETs. We hypothesize that privacy labels that take a contextual integrity approach—emphasizing which data flows are permitted and which are prohibited—could lead to improved comprehension of a variety of PETs (Nissenbaum, 2009). Our survey respondents found it more difficult to reason about the implications of local DP than central DP. This finding suggests that clearly explaining which data flows are permitted is particularly important for PETs that enable outsourced computation, such as local DP. Future work could confirm whether the techniques employed here, and the greater difficulty with mental model formation among participants, extends to other PETs that engage in outsourced computation, such as secure multi-party computation, trusted execution environments, and homomorphic encryption
@@ B.12. PETs
B.13. Background
How familiar are you with the following computer and Internet-related items? Please choose a number between 1 and 5, where 1 represents no understanding and 5 represents full understanding of the item. (Each item also offered ‘prefer not to answer’ option.)
@@ Appendix C Demographics
Table 6 describes the demographics of the 24 participants in the main interview study. Table 7 describes the demographics of the 10 participants who participated in the follow-up interviews. Table 8 summarizes the demographics of the survey respondents. Note that respondents could select multiple values for race/ethnicity and gender and that many respondents selected multiple options for race/ethnicity but did not explicitly describe themselves as multiracial. Table 9 displays the approximate percentage of respondents who expressed familiarity with various PETs out of all respondents who answered this question (n=684).
Table 6. Participant Demographics: Initial Interviews
@@
Table 7. Participant Demographics: Follow-up Interviews
@@
Table 8. Respondent Demographics
@@
Deliquescent security (distractor) 3 ¡1%
@@ Appendix D Designs
Table 10. Original Metaphor Descriptions
Figure 2. Top: Diagram for local model. Bottom: Diagram for central model.
@@
Figure 3. Original Privacy Labels.
Figure 4. Example of the Miro board setup used for the follow-up interviews.
Figure 5. In the initial interviews, we evaluated multiple versions of each explanation type for both the local and central models. Based on participant feedback, we dropped the diagram explanations and modified the privacy labels. During the follow-up interviews, we developed a new metaphor and introduced a text with information about the data protection process. Next, we compared the two privacy labels through a survey, and dropped the version with locks. Finally, we evaluated the disguise metaphor, the process text, and combinations of these texts and the privacy label with arrows.
Table 11. All Explanation Texts
@@ Appendix E Descriptive Statistics
Model Hack Law Org Graph Share
Metaphor
Central 0.78 0.05 0.54 0.14 0.89 0.03 0.41 0.46 0.54 0.22
Local 0.26 0.47 0.24 0.45 0.37 0.37 0.39 0.42 0.26 0.47
Process
Central 0.50 0.28 0.50 0.17 0.52 0.17 0.42 0.32 0.45 0.32
Local 0.28 0.48 0.3 0.50 0.38 0.42 0.32 0.28 0.28 0.52
Process+Metaphor
Central 0.79 0.11 0.76 0.05 0.92 0.08 0.61 0.34 0.50 0.37
Local 0.36 0.38 0.33 0.33 0.44 0.41 0.49 0.31 0.38 0.41
ArrowLabel
Central 0.92 0.03 0.95 0.00 0.92 0.05 0.58 0.26 0.82 0.00
Local 0.51 0.32 0.44 0.37 0.46 0.32 0.63 0.24 0.63 0.27
Label+Metaphor
Central 0.85 0.13 0.82 0.05 0.82 0.08 0.64 0.26 0.77 0.15
Local 0.72 0.22 0.58 0.33 0.58 0.33 0.69 0.19 0.61 0.31
Label+Process
Central 0.85 0.12 0.82 0.05 0.65 0.15 0.6 0.25 0.62 0.22
Local 0.59 0.28 0.44 0.31 0.67 0.21 0.62 0.18 0.67 0.28
Label+Process+Metaphor
Central 0.85 0.05 0.82 0.03 0.80 0.15 0.45 0.40 0.70 0.15
Local 0.56 0.26 0.59 0.18 0.56 0.26 0.69 0.15 0.49 0.26
Xiong
Central 0.91 0.03 0.44 0.12 0.62 0.15 0.35 0.44 0.44 0.32
Local 0.33 0.31 0.23 0.44 0.33 0.41 0.56 0.23 0.41 0.46
Table 12. Accuracy of Privacy Expectations

### v2

@@ 4.1.3 Data-Sharing Decision
5 Limitations
Our designs are limited in their focus on a single scenario. Although medical applications are often cited as motivation for studies of DP [48, 10], DP has not been widely deployed in medical contexts [20]. Nevertheless, our privacy labels are transportable to other domains. Future work could transfer our designs to other scenarios and test whether our findings still hold. A limitation of our evaluation is that encountering explanations of DP in practice differs significantly from encountering explanations in an online survey. Future work could investigate comprehension when these explanations are encountered in more natural settings. A third limitation is our focus on a U.S. audience. Our privacy labels may be received differently in a different cultural context. Finally, we present the nature of DP’s protection as binary, when in fact the level of protection depends on the choice of privacy budget. This simplification may be appropriate for small privacy budgets, but the question of determining an acceptable range for the privacy budget is itself a nontrivial problem.
@@ 5 Limitations
6 Discussion
Our results highlight the value of combining disparate best practices from prior work on explaining other security and privacy (S&P) concepts to explain complex PETs such as DP [53, 26]. We find that consequences-focused explanations (i.e., privacy label explanations that highlight information flows) to be a promising approach for promoting accurate understandings of potential data leaks in DP systems. However, to ensure that such explanations are trusted we find that it is necessary to pair such consequences-focused information with a limited amount of high-level information about mechanisms: how DP works to offer particular consequences and protections. Below we discuss potential pitfalls of privacy labels for DP as well as ways to extend our designs to explain other PETs individually or in combination.
@@ 6 Discussion
Although the nutrition label approach shows promise for setting appropriate privacy expectations, it is important to avoid pitfalls from prior deployments of nutrition labels for privacy [18]. For example, iOS privacy labels can be misleading and inaccurate [58], in part because developers struggle to create accurate labels [62]. Similarly, our labels for DP could be misleading if an organization has implemented DP incorrectly [46, 11, 15, 73, 65] or has chosen an inappropriately large privacy budget [27]. Specialized programming platforms, audits, and formal verification approaches are therefore an important complement to our work [69, 88, 107, 98, 25, 57], namely by ensuring that the communicated privacy guarantees match the implementation.
Furthermore, while privacy labels can empower individuals to make decisions that better align with their goals and values, it is also important not to overburden individuals in the same way that traditional privacy policies do [67]. As some of the participants we interviewed highlighted, it can be difficult to strike the right balance between simplicity and comprehensiveness. Such a balance is important not only for data subjects, but also for other audiences who may encounter DP. For instance, privacy labels for DP could be used to educate policymakers, advocacy organizations, or software developers to support them in various decision-making processes. For example, Mozilla’s “privacy not included” guide offers expert reviews to help buyers choose products that provide strong privacy and security, since it can be difficult for individual buyers to evaluate various data protection policies themselves. One could imagine a similar project to provide reviews for different data collection initiatives. An advocacy organization might use privacy labels for PETs like DP to identify and recommend certain initiatives that provide good S&P guarantees.
Finally, it is crucial that privacy labels for DP be contextual. While the information disclosures our explanations highlight are ones that people care about [19], they represent a starting point which should be used to further adapt explanations for specific contexts. The information disclosures we highlight may not be comprehensive of all specific disclosures people are concerned about across contexts. For example, privacy concerns in a particular educational setting may differ from a medical setting. Future work should also study ways to supplement privacy labels for DP with contextually-appropriate communication about the choice of privacy budget [77, 9].
@@
Our approach to designing privacy labels for DP could be adapted to other PETs. We hypothesize that privacy labels that take a contextual integrity approach—emphasizing which data flows are permitted and which are prohibited—could lead to improved comprehension of a variety of PETs [79]. Our survey respondents found it more difficult to reason about the implications of local DP than central DP. This finding suggests that clearly explaining which data flows are permitted is particularly important for PETs that enable outsourced computation, such as local DP. Future work could confirm whether the techniques employed here, and the greater difficulty with mental model formation among participants, extends to other PETs that engage in outsourced computation, such as secure multi-party computation, trusted execution environments, and homomorphic encryption
@@ B.12 PETs
B.13 Background
How familiar are you with the following computer and Internet-related items?
@@ Appendix C Demographics
Table C describes the demographics of the 24 participants in the main interview study. Table C describes the demographics of the 10 participants who participated in the follow-up interviews. Table C summarizes the demographics of the survey respondents. Note that respondents could select multiple values for race/ethnicity and gender and that many respondents selected multiple options for race/ethnicity but did not explicitly describe themselves as multiracial. Table C displays the approximate percentage of respondents who expressed familiarity with various PETs out of all respondents who answered this question (n=684).
Table 6: Participant Demographics: Initial Interviews
@@
Table 7: Participant Demographics: Follow-up Interviews
@@
Table 8: Respondent Demographics
@@
Deliquescent security (distractor) 3 <1%
@@ Appendix D Designs
Figure 2: Top: Diagram for local model. Bottom: Diagram for central model.
@@
Figure 3: Original Privacy Labels.
Figure 4: Example of the Miro board setup used for the follow-up interviews.
Figure 5: In the initial interviews, we evaluated multiple versions of each explanation type for both the local and central models. Based on participant feedback, we dropped the diagram explanations and modified the privacy labels. During the follow-up interviews, we developed a new metaphor and introduced a text with information about the data protection process. Next, we compared the two privacy labels through a survey, and dropped the version with locks. Finally, we evaluated the disguise metaphor, the process text, and combinations of these texts and the privacy label with arrows.
Table 11: All Explanation Texts
@@ Appendix E Descriptive Statistics
Table 12: Accuracy of Privacy Expectations
Model Explanation Hack Law Org Graph Share
Central Metaphor 0.78 0.05 0.54 0.14 0.89 0.03 0.41 0.46 0.54 0.22
Local Metaphor 0.26 0.47 0.24 0.45 0.37 0.37 0.39 0.42 0.26 0.47
Central Process 0.50 0.28 0.50 0.17 0.52 0.17 0.42 0.32 0.45 0.32
Local Process 0.28 0.48 0.3 0.50 0.38 0.42 0.32 0.28 0.28 0.52
Central Process+Metaphor 0.79 0.11 0.76 0.05 0.92 0.08 0.61 0.34 0.50 0.37
Local Process+Metaphor 0.36 0.38 0.33 0.33 0.44 0.41 0.49 0.31 0.38 0.41
Central ArrowLabel 0.92 0.03 0.95 0.00 0.92 0.05 0.58 0.26 0.82 0.00
Local ArrowLabel 0.51 0.32 0.44 0.37 0.46 0.32 0.63 0.24 0.63 0.27
Central Label+Metaphor 0.85 0.13 0.82 0.05 0.82 0.08 0.64 0.26 0.77 0.15
Local Label+Metaphor 0.72 0.22 0.58 0.33 0.58 0.33 0.69 0.19 0.61 0.31
Central Label+Process 0.85 0.12 0.82 0.05 0.65 0.15 0.6 0.25 0.62 0.22
Local Label+Process 0.59 0.28 0.44 0.31 0.67 0.21 0.62 0.18 0.67 0.28
Central Label+Process+Metaphor 0.85 0.05 0.82 0.03 0.80 0.15 0.45 0.40 0.70 0.15
Local Label+Process+Metaphor 0.56 0.26 0.59 0.18 0.56 0.26 0.69 0.15 0.49 0.26
Central Xiong 0.91 0.03 0.44 0.12 0.62 0.15 0.35 0.44 0.44 0.32
Local Xiong 0.33 0.31 0.23 0.44 0.33 0.41 0.56 0.23 0.41 0.46